# Exact black holes and black branes with bumpy horizons supported by superfluid pions

### v2

@@ I Introduction
Black holes and black branes are fundamental objects in both theoretical physics and astrophysics, whose properties are intrinsically determined by the shape of their horizons. In General Relativity (GR), in $3+1$ dimensions without cosmological constant, there are many rigidity theorems which restrict both the topology and the curvature of the horizon [33, 6, 28, 44, 22]. Now, the presence of a negative cosmological constant allows event horizons of constant curvature with toroidal or hyperbolic shapes [23, 40, 38, 1, 4, 47, 35], and in space-time dimensions higher than four, horizons with exotic topologies (such as black rings and black Saturns) can appear [15, 13] (see also Refs. [26, 39, 32]).
@@
A fundamental question therefore remains open: can bumpy horizons be supported in $3+1$ dimensions by a realistic matter field? This question is extremely relevant, for instance, in the analysis of the exterior gravitational field of compact objects with superfluid vortices piercing the external surface from the inside. Here, we answer this question in the affirmative. We consider GR with a cosmological constant coupled to a non-linear sigma model describing the low-energy dynamics of pions. Using an Ansatz for superfluid pion multi-vortices [5], the gravitational back reaction reduces to a Liouville equation with a smooth source that can be solved analytically. The matter content represents the pions superfluid, with the phase postulated by Feynman and Onsager [42, 20]. The resulting horizons are protected by an integer topological charge, the vorticity, which stabilizes the bumps against gravitational stretching.
@@ II The Einstein-NLSM
The Ansatz introduced in Ref. [5] is suitable to find the gravitational field corresponding to BPS multi-vortices:
@@
We take the following Ansatz for the metric
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}e^{P(x,y)}\left(dx^{2}+dy^{2}\right)\ .$ (4)
The above metric has been chosen searching for the most general static gravitational field with the property that the flat BPS equations for pion vortices imply the second order field equations.
The fact that $\partial_{t}$ is a Killing vector implies the conservation of the current $j^{\mu}=\sqrt{-g}\,T^{\mu}_{0}$. However, $j^{r}=0$, and therefore the integral of the energy density of the pionic field ($\varepsilon=j^{0}=\sqrt{-g}\,T^{t}_{\,\,t}$) over surfaces of constant $r$ is conserved. It may be written in a BPS form as
@@ IV.2 Effect of the vortices on the horizon
In general, the Liouville equation with sources (19) can only be integrated numerically. At the horizon $r=r_{H}$, this describes the deformation from spherical symmetry, the $u$ function, that the horizon gets due to the presence of matter through $H$ and $\Phi$. However, some of it behavior may be deduced qualitatively.
@@ IV.4 Mass and Entropy
The mass can be derived using the ADM formula. Because the metric has the exact same form than Schwarzschild or its dS or AdS counterparts, the only difference resides in the integral over the $\Sigma$ space, namely, the area $A$ in Eq. (20). In the $\gamma=1$ case, for instance, the mass is
@@
The entropy is also straightforward to obtain,
@@
The above expressions are typical of configurations with angular defects like global monopoles.
@@ V Final remarks
We have constructed the first analytic examples of black holes and black branes with bumpy horizons supported by realistic matter fields (pions) in GR minimally coupled with the NLSM. The number of bumps of the horizon is related to the quantized vorticity of the pions field, which exactly satisfies the Onsager-Feynman relation for the quantum superfluid phase. This discrete topological charge enters directly in the first law of black hole thermodynamics. This opens up remarkable new possibilities, such as the analysis of the gravitational field of compact objects where superfluid vortices pierce the surface of the object itself, which is extremely relevant in the description of pulsar glitches [43]. If realized in nature, such configurations would provide a concrete mechanism by which microscopic topological defects in a superfluid phase could imprint macroscopic structure on black hole horizons. Last but not least, our results allow an explicit realization of the holographic description of the entropy production and viscosity bounds due to broken translational symmetry.

### v3

@@ I Introduction
Black holes and black branes are fundamental objects in both theoretical physics and astrophysics, whose properties are intrinsically determined by the shape of their horizons. In General Relativity (GR), in $3+1$ dimensions without cosmological constant, there are many rigidity theorems which restrict both the topology and the curvature of the horizon [35, 8, 30, 46, 24]. Now, the presence of a negative cosmological constant allows event horizons of constant curvature with toroidal or hyperbolic shapes [25, 42, 40, 3, 6, 49, 37], and in space-time dimensions higher than four, horizons with exotic topologies (such as black ring and black Saturn) can appear [17, 15] (see also Refs. [28, 41, 34]).
@@
A fundamental question therefore remains open: can bumpy horizons be supported in $3+1$ dimensions by a realistic matter field? This question is extremely relevant, for instance, in the analysis of the exterior gravitational field of compact objects with superfluid vortices piercing the external surface from the inside. The evidence presented here confirms this to be the case. We consider GR with a cosmological constant coupled to a non-linear sigma model (NLSM) describing the low-energy dynamics of pions. Using an Ansatz for superfluid pion multi-vortices [7], the gravitational back reaction reduces to a Liouville equation with a smooth source that can be solved analytically. The matter content represents the pions superfluid, with the phase postulated by Feynman and Onsager [44, 22]. The resulting horizons are protected by an integer topological charge, the vorticity, which stabilizes the bumps against gravitational stretching.
@@ II The Einstein-NLSM
We take the following Ansatz for the metric
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}e^{P(x,y)}\left(dx^{2}+dy^{2}\right)\ ,$ (3)
where $(x,y)$ represent the coordinates on the transverse section, providing the base space for the pionic vortex configuration and the resulting horizon deformations discussed below.
For the matter fields we use the Ansatz introduced in Ref. [7], which is suitable to find the gravitational field corresponding to BPS multi-vortices:
@@
While our Ansatz may appear restrictive, it describes a configuration characterized by a single Killing vector associated with time translations; consequently, these solutions lack both spherical and axial symmetry. This reduction in symmetry is a direct consequence of the pionic vortices distributed at the horizon, where these topological solitons break rotational invariance and induce the characteristic deformations (bumps) of the geometry. Furthermore, as we will show below, despite that the matter field does not depend on the radial coordinate, the configuration possesses a finite mass and remains topologically protected.
The fact that $\partial_{t}$ is a Killing vector implies the conservation of the current $j^{\mu}=\sqrt{-g}\,T^{\mu}_{0}$. However, $j^{r}=0$, and therefore the integral of the energy density of the pionic field ($\varepsilon=j^{0}=\sqrt{-g}\,T^{t}_{\,\,t}$) over surfaces of constant $r$ is conserved.
The energy density is given by
$\varepsilon=\frac{K}{2\kappa}\left\{(\partial_{x}\alpha)^{2}+(\partial_{y}\alpha)^{2}+\sin^{2}\alpha\left((\partial_{x}\Phi)^{2}+(\partial_{y}\Phi)^{2}\right)\right\}\ ,$
and then, the energy may be written in a BPS form as
@@
It is a direct computation to show that all the solutions of the above BPS system are also solutions of the second order field equations:
$\Delta_{F}\Phi+2\cot(\alpha)(\vec{\nabla}\alpha\cdot\vec{\nabla}\Phi)=$ $0\ ,$ (8)
$\Delta_{F}\alpha-\frac{1}{2}\sin(2\alpha)(\vec{\nabla}\Phi)^{2}=$ $0\ ,$ (9)
being $\Delta_{F}=\partial_{x}^{2}+\partial_{y}^{2}$ the flat-space Laplacian. Note that in Eqs. (7) the metric does not appear, so that the above equations are exactly the same as in the flat space-time case [7].
@@ IV.1 Area of the horizon
The complete analysis of dynamical stability is beyond the scope of this manuscript. Nevertheless, the existence of a quantized topological charge, which appears explicitly through a discrete parameter in the solutions, is a highly desirable feature (absent in most of the available examples in the literature) that suggests intrinsic robustness and relevance of these solutions. We will come back on this issue in a future publication.
@@ IV.2 Effect of the vortices on the horizon
In general, the Liouville equation with sources (21) can only be integrated numerically. At the horizon $r=r_{H}$, this describes the deformation from spherical symmetry, the $u$ function, that the horizon gets due to the presence of matter through $H$ and $\Phi$. However, some of its behavior may be deduced qualitatively.
@@ IV.4 Mass and Entropy
The mass can be derived using the ADM formula [1]. Because the metric has the exact same form than Schwarzschild or its dS or AdS counterparts, the only difference resides in the integral over the $\Sigma$ space, namely, the area $A$ in Eq. (22). In the $\gamma=1$ case, for instance, the mass is
@@
The entropy is also straightforward to obtain [51],
@@
The above expressions are typical of configurations with angular defects like global monopoles [2].
@@ V Final remarks
We have constructed the first analytic examples of black holes and black branes with bumpy horizons supported by realistic matter fields (pions) in GR minimally coupled with the NLSM. The number of bumps of the horizon is related to the quantized vorticity of the pions field, which exactly satisfies the Onsager-Feynman relation for the quantum superfluid phase. This discrete topological charge enters directly in the first law of black hole thermodynamics. Furthermore, our results allow an explicit realization of the holographic description of the entropy production and viscosity bounds due to broken translational symmetry.
While astrophysical black holes likely involve diverse matter contributions, the role of pionic matter is uniquely significant due to its topological robustness. We have demonstrated that the resulting horizon bumps are associated with BPS topological vortices; since vorticity is topologically conserved, these bumps cannot be continuously turned off. This topological protection ensures that the signature of pionic superfluidity remains stable even when additional matter fields are present. Consequently, despite the complexities of realistic astrophysical environments, pionic vortices will manifest as persistent deformations of the black hole horizon.
An interesting outcome of the present paper is the following: bumps on the horizons are a fingerprint of superfluid pionic vortices. Such a fingerprint is topologically protected as vorticity is topologically conserved. Therefore, it is very interesting to search for such a fingerprint in astrophysical relevant situations where superfluid pionic vortices play a fundamental role (see, for instance, [45] and references therein). We will return to this issue in a future publication.